OpenStudy (anonymous):

Q: If a car accelerated from 5 m/s to 25 m/s in 10 seconds, how far will it travel?

OpenStudy (anonymous):

v = v0 + at 25 = 5 + 10a 25 - 5 = 10a 20 = 10a 20/10 = a a = 2m/s² v² = v0² + 2 * a * d (25² - 5²) / (2 * 2) = d (625 - 25) / 4 = d 600 / 2 = d d = 150m The car travel 150 meters :D

OpenStudy (kropot72):

A simple solution can be found as follows: Assuming the acceleration is uniform, the average speed during acceleration is \[\frac{5+25}{2}=15\ m/s\] The distance traveled in 10 seconds at an average speed of 15 m/s is given by: d = 15 * 10 = 150 m